# Look Twice Before You Answer: Memory-Space Visual Retracing for Hallucination Mitigation in Multimodal Large Language Models

### v1

@@ Abstract
Despite their impressive capabilities, Multimodal Large Language Models (MLLMs) are susceptible to hallucinations, especially assertively fabricating content not present in the visual inputs. To address the aforementioned challenge, we follow a common cognitive process - when one’s initial memory of critical on-sight details fades, it is intuitive to look at them a second time to seek a factual and accurate answer. Therefore, we introduce Memory-space Visual Retracing (MemVR), a novel hallucination mitigation paradigm that without the need for external knowledge retrieval or additional fine-tuning. In particular, we treat visual prompts as supplementary evidence to be reinjected into MLLMs via Feed Forward Network (FFN) as “key-value memory”, when the model is uncertain or even amnesic about question-relevant visual memories. Comprehensive experimental evaluations demonstrate that MemVR significantly mitigates hallucination issues across various MLLMs and excels in general benchmarks without incurring added time overhead, thus emphasizing its potential for widespread applicability.  [^1] Code can be available at https://github.com/1zhou-Wang/MemVR.
614
Figure 1: MemVR demonstrates strong performance across seven benchmarks spanning various domains (Left), with particularly outstanding results on the MME benchmark (Center). Additionally, MemVR is compared with contrastive decoding schemes, standing out for its ability to alleviate hallucinations using just a single inference, making it a more eco-friendly solution (Right).
@@ 1 Introduction
Multimodal Large Language Models (MLLMs), due to their formidable capacity to comprehend visual inputs, have emerged as indispensable tools in computer vision (Koh et al., 2024) and natural language processing (Tu et al., 2023) to tackle numerous visual tasks and facilitate complex visual question answering (VQA). Nevertheless, MLLMs still exhibit certain limitations, i.e., the so-called “hallucination” phenomenon (Huang et al., 2024b; Zheng et al., 2024). Specifically, MLLMs frequently generate descriptions inconsistent with user-provided visual inputs, such as incorrectly outputting non-existent things in the image or conflicting judgments. This flaw poses significant risks to the reliability of MLLMs as trustworthy assistants (Yu et al., 2024b), particularly in safety-critical applications (Zou et al., 2023) (e.g., clinical healthcare (Lin et al., 2024) and autonomous driving scenarios (Ding et al., 2024)). While the exact reasons for hallucinations in MLLMs are not fully understood, one possible factor could be the imbalance between their understanding of visual and textual information. Typically, MLLMs encode images into vision tokens via CLIP (Radford et al., 2021), which are fed along with text tokens into LLMs for decoding. LLMs excel at text comprehension but struggle with visual information perception and memory, where differences in information density between modalities may cause inconsistencies during decoding.
@@
Numerous methods have been proposed to mitigate hallucinations in MLLMs, and general studies can be broadly categorized into three streams: (i) Retrieval-Augmented Generation (RAG) (Shuster et al., 2021; Caffagni et al., 2024) that incorporates knowledge from external databases to mitigate the problem of “hallucination”, as well as (ii) through extra Fine-tuning (Yu et al., 2024b) to enhance the self-consistency of generation, and (iii) Contrastive Decoding (CD) strategy, which do not involve extra training. Specifically, RAG and fine-tuning patterns typically employ external knowledge retrieval or robust instruction-tuning datasets to post-hoc debias (Yang et al., 2024; Liu et al., 2023a), which inevitably introduces substantial computational overhead or storage requirements. For example, some approaches(Yin et al., 2023; Yu et al., 2024a) have fine-tuned models using high-quality visual instructions generated by advanced automated annotation tools including GPT-4 (OpenAI, 2023).
@@
CD-based methods (Li et al., 2023b; Shi et al., 2024) represent a simpler and more efficient way to mitigate hallucinations than RAG and fine-tuning based methods. Particularly, CD-based hallucination mitigation methods usually modulate the logits of the next token prediction through contrast manner or penalty mechanisms. As illustrated in Figure 2 (a), DoLa (Chuang et al., 2023) enriches factual knowledge via layer-wise contrasting and reduces the generation of incorrect facts in LLMs. In MLLMs, VCD (Leng et al., 2024) amplifies the language priors by adding Gaussian noise to the visual inputs, thereby reducing over-reliance on statistical biases and single-modal priors through contrasting output distributions from original and distorted visual inputs as in Figure 2 (b). This perturbation of original inputs requires task-specific design, inevitably doubling inference costs. More critically, contrastive distributions are agnostic to visual and instructional nuances, which may not always amplify the intended hallucinations, occasionally introducing potential noise into CD.
Figure 2: The conventional paradigm for hallucination mitigation, employing contrastive decoding methods such as DoLa (a) and VCD (b), is compared to our proposed MemVR strategy (c).
Method 20-Token Len 50-Token Len 80-Token Len
LLaVA-1.5 1880.3  $\downarrow$×$\times$1.0 3617.6  $\downarrow$×$\times$1.0 5256.6  $\downarrow$×$\times$1.0
+ VCD 4537.4  $\uparrow$×$\times$2.4 7690.8 $\uparrow$×$\times$2.1 11569.3 $\uparrow$×$\times$2.2
+ OPERA 6242.7  $\uparrow$×$\times$3.3 12672.3  $\uparrow$×$\times$3.5 19247.2  $\uparrow$×$\times$3.7
+ MemVR 1861.7  $\uparrow$×$\times$1.0 4000.9  $\uparrow$×$\times$1.1 5545.5  $\uparrow$×$\times$1.1
Table 1: Efficiency Comparisons for generating different length tokens, using an NVIDIA-A40 GPU. Inference time (ms) of different methods is recorded.
@@
In this work, we delve into the challenges of hallucination mitigation in MLLMs and address the shortcomings of CD-based approaches. Our research is grounded in a common cognitive process: when the initial memory of certain critical visual details fades, it is intuitive to look at them for the second time to search for the accurate answer (O’regan, 1992; Ballard et al., 1995; Horowitz & Wolfe, 1998). Different from visual contrastive decoding strategies that alleviate hallucinations by diminishing language priors (Leng et al., 2024; Qu et al., 2024; Park et al., 2024), we propose a novel Memory-space Visual Retracing (MemVR) method that mitigates hallucinations through supplementing visual evidence, akin to the two sides of a coin. MemVR, as shown in Figure 2 (c), is an architecture-agnostic, plug-and-play solution that re-injects visual features into an intermediate layer suffering from vision-related memory lapse with only one regular inference. This novel hallucination mitigation paradigm in terms of efficiency, and its inference cost and performance are optimal compared to previous studies as listed Table 1. It’s a game-changer for effectiveness and efficiency. Through extensive experiments on multimodal hallucination benchmarks, as well as GPT-4o [^2] GPT-4o-2024-08-06: https://platform.openai.com/docs/models/gpt-4o. evaluations, we show the comprehensive performance improvements of MemVR in hallucination mitigation and general capabilities. Our contributions can be summarized as follows:
  • We propose MemVR, a novel training-free hallucination mitigation paradigm that effectively alleviates hallucinations in MLLMs. In contrast to previous methods, which primarily focus on eliminating biases of language priors, MemVR seeks to replenish question-relevant visual clues towards more evidential responses, which signifies the other side of the coin.
  • We design a dynamic premature layer injection strategy with visual retracing in MLLMs, mimicking human intuitive thinking to revisit image features for self-consistency and credible answers when pivotal memories are scrambled. Furthermore, we theoretically demonstrate that visual retracing can effectively diminish hallucinations from an information-theoretic perspective.
  • Comprehensive experimental results demonstrate the effectiveness of MemVR in mitigating hallucinations and enhancing general cognitive and perceptual performance, as well as its high efficiency. Our research will make a substantial contribution to trustworthy multimodal intelligence.
@@ 2 Related Work
MLLMs and Challenges. In recent years, MLLMs have made remarkable progress, particularly as they have evolved from the foundations laid by Vision Language Models (VLMs). Early based on BERT-style language decoders (Devlin, 2018), which achieved initial cross-modal integration by combining visual and textual data (Li et al., 2022). Leveraging open-source Large Language Models (LLMs) such as LLaMA families (Touvron et al., 2023), MLLMs (Alayrac et al., 2022; Wu et al., 2024) have demonstrated enhanced adaptability across a range of visual language tasks, leading to a more profound ability to interpret the world. Models like LLaVA (Liu et al., 2024), Qwen-VL (Bai et al., 2023), and GLM4V (Wang et al., 2023) have further advanced this field, enabling users to interact with these agents using both image and text prompts. These models adhere to two critical training phases: pre-training feature alignment and instruction fine-tuning, ensuring they better comprehend the format of instruction inputs (Yin et al., 2024). However, despite their impressive performance in many areas, MLLMs still suffer from hallucination issues. Thus, in this paper, we primarily conducted experiments and analysis on these three representative models.
Hallucinations in MLLMs. Before LLMs, hallucination in NLP was mainly seen as generating nonsensical or deviant content. In MLLMs, “hallucination” is defined as the model generating content inconsistent with the provided image. This issue stems significantly from inadequate alignment among modalities. Several methods have been explored to mitigate hallucinations in MLLMs. Early efforts focused on fine-grained modality alignment (Rohrbach et al., 2018) and reducing co-occurrence biases (Kim et al., 2023) in small-scale VLMs, but these approaches struggle to scale with MLLMs. More recent strategies involve hallucination-targeted datasets for fine-tuning (Gunjal et al., 2024), post-hoc revisors (Zhou et al., 2024), and adopting RLHF (Yu et al., 2024b). While effective, these methods are resource-intensive. CD-based approaches (Chuang et al., 2023; Leng et al., 2024) adjust the decoding distribution to mitigate hallucinations, but it does not consistently improve performance. Compared with them, our MemVR stands as “a paradigm of effectiveness and efficiency” in hallucination mitigation, effortlessly enhancing performance without extra training.

### v2

@@ Abstract
Despite their impressive capabilities, multimodal large language models (MLLMs) are prone to hallucinations, i.e., the generated content that is nonsensical or unfaithful to input sources. Unlike in LLMs, hallucinations in MLLMs often stem from the sensitivity of text decoder to visual tokens, leading to a phenomenon akin to “amnesia” about visual information. To address this issue, we propose MemVR, a novel decoding paradigm inspired by common cognition: when the memory of an image seen the moment before is forgotten, people will look at it again for factual answers. Following this principle, we treat visual tokens as supplementary evidence, re-injecting them into the MLLM through Feed Forward Network (FFN) as “key-value memory” at the middle trigger layer. This look-twice mechanism occurs when the model exhibits high uncertainty during inference, effectively enhancing factual alignment. Comprehensive experimental evaluations demonstrate that MemVR significantly mitigates hallucination across various MLLMs and excels in general benchmarks without incurring additional time overhead. The implementation is available from https://github.com/1zhou-Wang/MemVR.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs), known for their ability to process visual, auditory and textual data, are crucial in fields such as computer vision (Koh et al., 2024) and natural language processing (Tu et al., 2023), helping with visual tasks and complex visual question answering. However, MLLMs still face challenges, notably the “hallucination” issue (Huang et al., 2024b; Zheng et al., 2024; Lyu et al., 2025), where they generate contents inconsistent with original inputs, such as generating nonexistent objects or conflicting judgments. This flaw undermines their reliability, especially in areas critical to safety such as healthcare (Lin et al., 2024) and autonomous driving (Ding et al., 2024). Although the causes of hallucinations are unclear, one potential factor is the imbalance between their understanding of visual and textual modalities. This imbalance may induce biases when the model integrates multimodal information, leading to outputs that do not match objective facts.
@@
Currently, numerous methods are being tried out to solve this problem. General studies can be broadly categorized into four streams: (i) Retrieval-Augmented Generation (RAG) (Qu et al., 2024) which incorporates knowledge from external databases to mitigate hallucinations, as well as (ii) through extra fine-tuning (Yu et al., 2024a) to enhance the self-consistency of generation; (iii) attention intervention (Huang et al., 2024a) and (iv) Contrastive Decoding (CD) (Leng et al., 2024) strategies, which not involve extra training. Specifically, RAG and fine-tuning patterns typically employ external knowledge retrieval or robust instruction-tuning datasets to post-hoc debias (Yang et al., 2024; Liu et al., 2023a), which inevitably introduces substantial computational overhead or storage requirements. Attention intervention, though not requiring additional data, usually involves retrospection-allocation operations, which bring about high inference latency and a large memory footprint.
Figure 1: Comparison of the conventional CD-based hallucination mitigation paradigm VCD, and our proposed efficient MemVR.
@@
CD-based methods (Li et al., 2023a; Shi et al., 2024) represent a simpler and more efficient way to mitigate hallucinations than other paradigms. Specifically, CD-based hallucination mitigation paradigms, represented by VCD (Leng et al., 2024), modulate logits of the next token prediction in a contrastive manner. As shown in Figure 1 left, VCD amplifies the language priors by adding Gaussian noise to the visual inputs, reducing over-reliance on statistical biases and single-modal priors through contrasting output distributions from original and distorted visual inputs. This perturbation of original inputs requires task-specific design, inevitably ✄ doubling inference costs. More critically, contrastive distributions are agnostic to visual and instructional nuances, which may not always amplify the intended hallucinations, occasionally ✄ introducing potential noise into CD.
@@
In this work, we delve into the challenges of hallucination mitigation in MLLMs and propose a novel paradigm that overcomes the two limitations of CD-based approaches. Our research is grounded in a common cognitive process: when the memory of an image seen the moment before is forgotten, it is intuitive to look twice for accurate and factual answers (Ballard et al., 1995; Horowitz & Wolfe, 1998). Following this principle, we propose Memory-space Visual Retracing (MemVR) that mitigates hallucinations through supplementing visual evidence, which can also be called look-twice mechanism. As shown in Figure 1 right, MemVR reinjects visual tokens through Visual Retracing (VR), i.e., look-twice mechanism, into the middle trigger layer suffering from high uncertainty, without incurring additional inference cost. Compared with VCD and other approaches, our proposed look-twice mechanism is optimal in terms of performance, efficiency, and memory cost as shown in Figure 2 and Table 1. Through extensive experiments on multimodal hallucination benchmarks, as well as GPT-4o evaluations, including eight public benchmarks, we show the comprehensive performance improvements of MemVR in hallucination mitigation and general capabilities. The main contributions can be summarized as follows:
  • We propose MemVR, a novel, efficient, minimalist, and plug-and-play approach that achieves both model fidelity and efficiency, which reinforces attention to visual information for enhancing modality balance during the forward pass, without eliminating beneficial language priors.
  • We present static and dynamic VR strategies that shift hidden states of the intermediate layer in MLLM for self-enhancement, rather than modulating logits directly in a CD manner, thus avoiding multi-round decoding.
  • Our analysis reveals that hallucinations are triggered by the sensitivity of text decoder (i.e., LLM) to non-text modality. This finding is experimentally validated.
  • Comprehensive experiments and evaluations on multiple models demonstrate that MemVR outperforms SOTA methods in performance and inference speed, such as +7.0% on the POPE benchmark, 15.6% improvement on CHAIRI, and total score of MME increased by +32.2 marks. The results on popular hallucination and general benchmarks validate the generalizability of our method.
Figure 2: Radar charts comparing models across benchmarks.
Method Latency $\downarrow$  (ms/token)    Throughput $\uparrow$   (token/ms) Time Cost $\downarrow$ for 80 Tokens (ms) Memory $\downarrow$ Cost (MB)
Greedy 65.71 ($\times$1.00) 0.015 ($\times$1.00) 5256.6 ($\times$1.00) 14257 ($\times$1.00)
Sample 69.84 ($\times$1.06) 0.015 ($\times$1.00) 5587.0 ($\times$1.00) 14262 ($\times$1.00)
OPERA 240.59 ($\times$3.66) 0.004 ($\times$0.27) 19247.2 ($\times$3.66) 21300 ($\times$1.49)
ICD 70.84 ($\times$1.08) 0.014 ($\times$0.93) 5666.9 ($\times$1.08) 14263 ($\times$1.00)
VCD 144.62 ($\times$2.20) 0.007 ($\times$0.47) 11569.3 ($\times$2.20) 14967 ($\times$1.05)
MemVR 68.32 ($\times$1.04) 0.015 ($\times$1.00) 5545.5 ($\times$1.06) 14345 ($\times$1.01)
Table 1: Performance comparison of SOTA methods and MemVR in latency, throughput, time cost, and memory usage. The best and suboptimal results are highlighted in green and blue, respectively.
@@ 2 Related Work
Table 2: Comprehensive comparisons between the proposed MemVR and existing approaches are presented. MemVR introduces a low-latency look-twice decoding mechanism, optimizing hidden states to support multimodal integration and enhance overall performance. MemVR uniquely achieves visual hallucination mitigation and general improvement. SOTA methods we compare are emphasized in gray.
Methods
 VH
   Mitigation
General
Improvement
Inference
Latency
Expand to
More Modalities
Modified
Component(s)
Decoding Paradigm
DoLa (Chuang et al., 2023) Negative Low Logits Contrastive
OPERA (Huang et al., 2024a) Medium High Attention matrix Att-intervention
EAH (Zhang et al., 2024) Medium High Attention matrix Att-intervention
CCA (Xing et al., 2024) Medium High Attention matrix Att-intervention
ICD (Wang et al., 2024) Medium Medium Textual input, logits Contrastive
ID (Kim et al., 2024) Medium Medium Textual input, logits Contrastive
SID (Huo et al., 2024) Medium Medium Visual tokens, logits Contrastive
VCD (Leng et al., 2024) Low Medium Visual input, logits Contrastive
HALC (Chen et al., 2024b) Medium Ultrahigh Visual input, logits Contrastive
VORD (Neo & Chen, 2024) Medium Medium Visual input, logits Contrastive
MemVR (ours) High Low Hidden states Look-twice
@@ Methods
MLLMs and Challenges. In recent years, MLLMs have made remarkable progress, particularly as they have evolved from the foundations laid by Vision Language Models (VLMs). Early based on BERT-style language decoders (Devlin, 2018), which achieved initial cross-modal integration by combining visual and textual data (Li et al., 2022). Leveraging open-source Large Language Models (LLMs) such as LLaMA families (Touvron et al., 2023), MLLMs (Alayrac et al., 2022; Wu et al., 2024) have demonstrated enhanced adaptability across a range of visual language tasks, leading to a more profound ability to interpret the world. Models like LLaVA (Liu et al., 2024b), Qwen-VL (Bai et al., 2023), GLM4V (Wang et al., 2023), and LLaVA-Next (Liu et al., 2024a) have further advanced this field, enabling users to interact with these agents using both image and text prompts. These models adhere to two critical training phases: pre-training feature alignment and instruction fine-tuning, ensuring they better comprehend the format of instruction inputs (Yin et al., 2024). However, despite their impressive performance in many areas, multimodal large language models still suffer from hallucination issues. In this work, we conducted experiments and analysis on these representative models to validate MemVR’s applicability.
Mitigating Hallucinations in MLLMs. Researchers have made extensive efforts to uncover the causes of hallucinations (Yin et al., 2023; Zhou et al., 2023; Bai et al., 2024). Early works to mitigate hallucinations focused on fine-grained modality alignment (Rohrbach et al., 2018) and reducing co-occurrence biases (Kim et al., 2023) in small-scale models. More recent strategies involve hallucination-related datasets for fine-tuning (Gunjal et al., 2024), post-hoc revisors (Zhou et al., 2024), and adopting RLHF (Yu et al., 2024a). LURE (Zhou et al., 2023) trains a reviser to edit the possible hallucinated words in the responses. While effective, these methods are resource-intensive. Attentional intervention strategies (Zhang et al., 2024; Xing et al., 2024), represented by OPERA (Huang et al., 2024a), are simpler and do not require additional data or training, but have a higher inference latency. CD-based approaches (Chuang et al., 2023; Chen et al., 2024b; Neo & Chen, 2024), represented by ICD (Wang et al., 2024) and VCD (Leng et al., 2024), adjust the decoding distribution to mitigate hallucinations in MLLM, but this does not consistently improve performance as it introduces potential noise into the CD.
Comparisons. Distinct from these methods, our proposed MemVR offers a novel decoding strategy that effectively reduces visual hallucinations (VH) without necessitating extra models, data, and training. Table 2 illustrates the differences and advantages of our method compared to recent representative SOTA approaches. DoLa (Chuang et al., 2023) targets hallucinations in LLM and is negative for alleviating VH, although it exhibits low latency. Attentional intervention methods (Zhang et al., 2024; Xing et al., 2024), represented by OPERA (Huang et al., 2024a), have achieved considerable progress in mitigating VH, but are limited to the problem of high latency. CD-based approaches, whether modifying the textual input (Wang et al., 2024; Kim et al., 2024) or visual input (Leng et al., 2024; Chen et al., 2024b; Neo & Chen, 2024), both aim to reduce the output probability of incorrect tokens through the comparison of logits. However, this brings two challenges, one is that the CD strategy may introduce noise to the output distribution, thus losing the original capability, and the other is that the CD-based mechanisms often require multiple rounds of inference to obtain several pairs of logits for contrasting, resulting in a high latency. Significantly, these methods work negatively or fail in general-purpose testing. Compared with them, MemVR stands as “a paradigm of effectiveness and efficiency” in visual hallucination mitigation and general improvement.
3 Background and Motivation
3.1 Problem Formulation
Given an MLLM $\mathcal{M}_{\theta}^{\text{MLLM}}$ parameterized by $\theta$, with a general architecture consisting of a text embedding layer, a vision encoder, a vision-text interface module, a text decoder consisting of $L$ number of transformer layers, and an affine layer $\varsigma(\cdot)$ which predicts the distribution of the next token. For an image-grounded text generation task, given a textual query $x$ and an input image $v$, MLLM first extracts vision features of $v$ by the vision encoder, and then converts them into visual tokens $z_{v}$ by MLP or Q-Former (Wadekar et al., 2024) modules. Aligned vision tokens $z_{v}$ are concatenated with the query $x$ as input to the text decoder, and finally decoded into a textual response $y$ autoregressive, which is formulated as follows:
$y_{t}\sim p_{\theta}(\cdot|v,x,y_{<t})\propto\text{softmax}(f_{\theta}(\cdot|v,x,y_{<t})),$ (1)
where $y_{t}$ indicates the $t^{th}$ token, $y_{<t}$ is the token sequence generated up to time step $t$, and $f_{\theta}$ is the logit distribution, i.e., unnormalized log-probabilities produced by $\mathcal{M}_{\theta}^{\text{MLLM}}$.
When the text generation $y$ is inconsistent or in conflict with the input image $v$, MLLM is believed to present hallucination issues. The objective of visual hallucination mitigation is to minimize the appearance of incorrect or conflicting tokens, ensure faithfulness to $v$ when answering the query $x$, and simultaneously maintain high-quality text generations.
3.2 Why Does MLLM Generate Hallucination?
Hallucinations in MLLM are caused by multiple factors, including inherent biases in the training data (Zhou et al., 2023), visual uncertainty resulting from the model’s statistical bias and priors (Leng et al., 2023), and the limitations of current models in accurately discerning context and fact throughout the output generation process (Daunhawer et al., 2021). Upon a more in-depth analysis, we consider that the imbalance of modalities in MLLM and the autoregressive characteristic of language models are likely crucial factors causing their hallucinatory phenomena. Taking image and text as an example, since an image possesses a much higher information density than a piece of text, it is reasonable to suppose that LLMs struggle to understand or memorize vision information compared to text. Moreover, autoregressive decoding causes MLLMs to increasingly depend on textual information, including query $x$ and growing tokens $y_{<t}$, inevitably decreasing reliance on visual input. As PAI(Liu et al., 2024c) proposed to mitigate hallucinations by paying more attention to images, this reflects the fact that hallucinations may be caused by modality imbalances.
Figure 3: Uncertainty of different layers to predict the next token. Rows denote indices of the early layers, and column names are decoded tokens in each step. Uncertainty distribution is dynamic.
Figure 4: (Left) Performance under different scaling ratios to text / image feature value on MME; (Right) Performance changes when look-twice to text / image / text+image (i.e, @t+i), respectively.
To further verify this conjecture, we scale the feature values of different modalities up or down to simulate the modality imbalance phenomenon. As shown in the left of Figure 4, it can be observed that: when the image and text features are scaled up proportionally, modality balance is disrupted in both cases, yet the performance decline caused by magnifying the image features is more significant than that of text features; however, when the image and text features is scaled down proportionally, the contraction on text features led to a drastic performance drop, while the scaling down of visual features had a relatively minor impact. We can obtain three findings in this experiment: ① modality imbalance issues cause hallucinations in MLLMs, ② text decoders (i.e., LLMs) are more text-informed, ③ LLMs are harder to comprehend the visual modality than textual inputs.
This modality imbalance leads to a substantial deviation from the accurate representation of visual input, eventually giving rise to hallucinations, as manifested by the phenomenons in the aforementioned studies (Zhou et al., 2023). This is especially evident in generating longer responses, explaining the correlation between higher VH and larger maximum token lengths, as presented in Huang et al. (2023).
Figure 5: Uncertainty distribution across layers during token reasoning in hallucinations. Red-outlined regions show higher uncertainty in middle and late layers for hallucinatory tokens: a, pom.
3.3 The Pattern of Hallucinations in MLLM
In order to further explore the potential pattern of hallucinations in MLLM, this study employs uncertainty as the metric. As findings of Chen et al. (2024a) in LLMs: incorrect tokens generally exhibit higher entropy than correct ones, we also observe this phenomenon in MLLMs, the visualization case is shown in Figure 5. From Figure 5, it can be noticed that when the model generates illusion tokens, i.e., a, pom -eg-ran-ate (pomegranate), the uncertainty in the middle and last layers of the model is high in MLLM.
Uncertainty quantification. Following the DoLa (Chuang et al., 2023), we compute the probability of the next token via the vocabulary head $\varsigma(\cdot)$ on each layer during reasoning. Then, we introduce an entropy-based metric (Farquhar et al., 2024) to quantify the output uncertainty of each layer in the text decoder as $u=\sum-p_{i}\log p_{i}/\log{N}$, where $\{p_{i}\}_{i=1}^{N}$.
In addition, in the context of tokens involving objects, attributes or relations, uncertainty is also high. We conduct a preliminary analysis with 32-layer LLaVA-1.5-7B. Specifically, we compute the uncertainty in the output distributions of early layers. Figure 3 shows the uncertainty scores of different early layers when decoding the answer, we can observe that the computed uncertainty remains relatively high in later layers when predicting key entity objects, attributes, or relations, such as wooden, filled, foods. This phenomenon suggests that LLM is still uncertain about its predictions in the last few layers and may inject more factual knowledge into the predictions. On the other hand, when predicting function words and those tokens copied from the question, e.g., image, with, we observe that the uncertainty becomes very low from the middle layers. This finding implies that the model is deterministic for easy-to-predict tokens at the intermediate layer and keeps the distribution of outputs almost constant at higher layers, however, it is more uncertain for difficult-to-predict key tokens and may constantly change its predictions until the final layer.
3.4 Refresh Visual Memory Alleviates Hallucination
Based on the findings, we propose that both overreliance on textual information and schema imbalance contribute to a phenomenon akin to “amnesia” about visual information. It is easy to understand that from the shallow to the deep layers of the Transformer, attention is progressively biased towards textual tokens, resulting in visual tokens at the deeper layer that hardly affect the results, which is consistent with how the attention intervention strategies work. To address this, we try to refresh visual information for trustworthy answers when the model encounters high uncertainty.
We verify this hypothesis through an empirical study. As shown in the right of Figure 4, the proposed look-twice strategy is used for combinations of different modalities, including text, image, text+image, where it can be found that the best performance improvement is achieved while only replenishing image information to the model, i.e., refreshing visual memory can effectively alleviate hallucinations in MLLMs. This study lends further credence to our conjecture regarding the causes of hallucinations in Sec. §3.2.